# Population Protocols Revisited: Parity and Beyond

### v2

@@ 2 The Parity Problem
Recall that in the parity problem, agents are tasked with computing the parity of a given sub-population size. In this work, we present a stable population protocol that solves the parity computation problem in polylogarithmic time. A common strategy for designing such protocols involves running a fast Monte Carlo algorithm that solves the problem with high probability, complemented by a slower, stable backup when needed. This approach has proven effective for problems such as majority computation and leader election, as discussed in the introduction. We adopt the same methodology in our construction. Notably, our protocol ensures both stabilisation and silence.
We begin (Section 2.1) by presenting a simple protocol that computes the parity of a given subpopulation $X$ of agents, using a fixed number of states. It always stabilises in time $O(n\log n)$, both in expectation and with high probability. To our knowledge, it represents the current state-of-the-art for stable parity computation and serves as the backup in our overall protocol design.
@@ 2.1 Slow Stable Parity Protocol
The slow protocol builds on the leader election protocol of [5]. It uses four states: $L_{0},L_{1},F_{0},F_{1}$}, with $L_{x}$ denoting leaders and $F_{x}$ followers. Initially, all agents in subpopulation $X$ start in $L_{1}$; all others start in $F_{0}$. On the interaction of two leaders, the rule $L_{x}+L_{y}\rightarrow L_{(x+y)\bmod 2}+F_{0}$ applies, where one remains a leader (with updated parity), and the other becomes a follower. Eventually, a single leader remains, encoding the parity of $|X|$ in its state index. This is then propagated via rule $L_{x}+F_{y}\rightarrow L_{x}+F_{x}$, so all followers adopt the leader’s index.
@@
2.2 Fast Parity Protocol
In this section, we provide an outline of the first fast, stable, and silent parity protocol (Algorithm 1).
@@ 2.2 Fast Parity Protocol
The balance scale idea stems from recomputation of powers of 2 (due to limited space), as proposed in [6]. Equipped with polylogarithmic memory we can afford initial construction of subpopulations $W_{i}$ of cardinality $2^{i},$ for all $i=0,1,\dots,\lfloor\log\frac{n}{3}\rfloor$. These subpopulations serve as weights in balance scales (Figure 2) implemented via the fast majority protocol. Using subpopulation $W_{i}$, we can determine whether $|X|-|Y|\geq 2^{i}$ by computing exact majority (with tie detection) between $|X|$ and $|Y|+|W_{i}|$. This balancing mechanism enables efficient computation of parity and size of any subpopulation.
@@
There exists a population protocol which solves the parity problem utilising $O(\log^{3}n)$ states and stabilises silently in expected $O(\log^{3}n)$ time.
The proof of Theorem 1 is provided in Section 3.6, near the end of the main part of this paper.
@@ 3 Efficient Protocols with Anomaly Detection
In this section, we develop methods for constructing efficient population protocols that silently stabilise with certainty. Specifically, assuming the existence of two protocols: a slow, state-efficient protocol that stabilises silently in polynomial time with probability 1, and a fast Monte Carlo type protocol possibly composed of subprotocols, we show how to integrate these two to obtain a protocol that is both efficient in time and guaranteed to stabilise silently.
Monte Carlo protocols typically aim to compute the desired solution with high probability, but the correctness of the solution remains uncertain after execution. Existing mechanisms for detecting fast protocol failures and transitioning to a consistently stabilising protocol have been developed on an ad hoc basis. In contrast, we propose a novel universal anomaly-detection mechanism that effectively integrates multiple Monte Carlo subprotocols with a reliable, though slower, always-correct solution. Specifically, this mechanism identifies efficiently scenarios necessitating a shift to a slower, always-stabilising protocol and facilitates the creation of complex population protocols by sequentially combining simpler ones. Additionally, the constructed protocols are silent.
Our efficient (polylogarithmic time and space) protocols converge to the correct solution with high probability or signal potential errors via an efficient anomaly-detection mechanism. Such signal is treated as a failure of the fast protocol, even if it does not always indicate an actual error. Our universal strategy is to first run a fast Monte Carlo protocol that either stabilises with high probability or, upon detecting an anomaly, switches to a slower, always-stabilising fallback.
@@ 3.1 Clock Protocol
Central to our design is the clock protocol, which acts as a clocking device and a command counter, tracking algorithmic progress and synchronizing all agents to the same point in time. In rare cases of desynchronisation, it signals an anomaly. Each main (e.g., parity) protocol is built by sequentially composing subprotocols, each spanning one or more rounds. The clock counts down $R=\text{poly}\log(n)$ rounds of the main protocol, each taking $O(\log n)$ time. To measure a round’s duration, a leader-maintained interaction counter tracks $c\log n$ interactions, triggering the transition to the next round.
@@ 3.2 MC+ Protocols
In this work, we focus on fast population protocols, specifically Monte Carlo protocols with anomaly detection (MC+ protocols). These protocols either silently stabilise on the correct computational output whp or trigger an anomaly detection signal. To maintain order in protocol analysis, we distinguish two types of anomalies:
@@
If a fast MC+ protocol and a polynomial-time stable protocol exist for the same problem, they can be combined to yield a protocol that is both fast (in expected running time) and stable (with high probability). Moreover, if the polynomial-time stable protocol is silent, the resulting combined protocol is also silent.
@@
Assume that the fast protocol completes its computations in $R$ rounds. If not successful, it reports a protocol anomaly in at most $n^{a}$ average time with negligible probability less than $n^{-a}$, for suitably chosen protocol parameters and $a>0$. In this time, by Lemma 3 the clock reports an anomaly with negligible probability less than $n^{-a}$. Assume also that the slow algorithm runs in expected time less than $n^{a}$. The expected execution time of the combined algorithms consists of the following components:
@@
Therefore, the expected execution time of the combination of fast and slow protocol is at most $O(R\log n)+1+2=O(R\log n).$ Moreover, the combination of the protocols is stable and silent.
@@
3.3 Leader Election
The leader election protocol proposed in this work builds upon a scheme that has been employed in several prior studies [6, 8, 13]. In this approach, the leaders perform coin flips in successive rounds. Whether an agent draws heads or tails depends on the random scheduler’s designation of that agent as either the initiator or the responder in an interaction. If, in a given round, one leader flips tails while another flips heads, the leader obtaining tails becomes a follower. Unlike previous protocols, our method associates the coin-flip outcome (heads or tails) with the round number, thereby ensuring that at least one leader always persists. Apart from clock anomalies that can be detected during the execution of the protocol, the only other potential failure arises when multiple leaders remain. This is identified as a protocol anomaly when two such leaders interact after the leader election has concluded. This detection occurs in the expected time $O(n)$.
@@ 3.3 Leader Election
3.4 Exact Majority
We now construct and analyse our majority computation protocol. This protocol operates over $LOG+4$ rounds and achieves a time complexity of $O(\log^{2}n)$. Throughout this construction, we assume that a unique leader has already been selected. The protocol compares two subpopulations, $X$ and $Y$, determining which is larger or whether they are of equal size. Notably, $X$ and $Y$ need not be disjoint. At the start of the first round, agents are assigned states as follows: elements exclusive to $X$ are assigned state $(1)$; elements exclusive to $Y$ state (1)(-1); elements in both $X$ and $Y$ or in neither state $(0)$. During each round $r$, an averaging operation is performed on pairs of agents in round $r$: $(a)+(b)\rightarrow\left(\lfloor\frac{a+b}{2}\rfloor\right)+\left(\lceil\frac{a+b}{2}\rceil\right)$. When transitioning to the next round, a doubling operation is applied: $(a)\rightarrow(2a)$.
@@ 3.4 Exact Majority
Suppose we are given an initial configuration in which each agent has a state $(a)$, where $a$ is an integer and the discrepancy between states is $\delta$. A protocol using only averaging arrives at a configuration with a discrepancy between states of at most $2$ after time $O(\log\delta+\log n)$ whp.
@@
Let bias $\Delta=|X|-|Y|$ and ${\cal S}_{r}$ be the sum of states $(a)$ in which each agent starts round $r$. Note that ${\cal S}_{1}=\Delta$. The averaging operation does not change the sum of agent states $(a)$ in round $r$. Thus as long as there are no $(x_{win})$ or $(y_{win})$ states, ${\cal S}_{r}$ is also the sum of states $(a)$ in which each agent ends round $r$. Due to doubling operation ${\cal S}_{r+1}=2{\cal S}_{r}$, as long as states $(x_{win})$ and $(y_{win})$ are not present.
@@
3.5 System of Weights
In this section, we formulate the MC+ protocol, which creates weights that are disjoint sets of agents. Weight $W_{i}$ is a set consisting of $2^{i}$ agents. The weight creation rounds are indexed by integers $r=0,1,2,\ldots,LOG+2$. In round $r$, the protocol creates $W_{r}$, consisting of $2^{r}$ agents. If each of these agents is assigned unit mass, then $W_{r}$, composed of $2^{r}$ agents, has a total mass of $2^{r}$. At the start of the protocol, all followers are in state $\phi$, which marks free agents. At the beginning of each subsequent round $r$, if the leader has not yet reached a final state $max(r)$ or $Max(r)$, it adopts state $\lambda_{r}$, which serves as the precursor of $W_{r}$ carrying its mass $2^{r}$. The leader then transfers this mass to the first free agent it encounters via transition: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$, where $\lambda$ denotes a non-final leader state that carries no mass. Agents participating in the creation of $W_{r}$ have states $w_{r}(i)$, indicating that the agent possessing it belongs to $W_{r}$ and holds mass $2^{i}$ from that weight. Creating the weight is accomplished by a transition that divides the mass greater than $2^{0}$ between two agents: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$. This process completes when all agents in $W_{r}$ have mass $1$, that is, when they reach state $w_{r}(0)$.
@@ 3.5 System of Weights
4:  The leader adopts state $\lambda$ and non-leaders adopt state $\phi$.
5:  for round $r\leftarrow 0$ to $LOG+2$ do
6:   Upon entering round $r$
7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $2^{r}$).
8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$.
9:   The leader transfers mass $2^{r}$ to the first free agent it encounters:
10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$
11:   Distribute the mass by repeating: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$ for $i>0$
12:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
13:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
14:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
15:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
16:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
17:      it adopts state Max(r2)Max(r-2).
18:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
19:      it marks state $w_{r-3}$ as top.
20:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
21:   if a non-leader in state max(r2)max(r-2) enters round $r$
22:       interacting with state $\neq$ Max(r2)Max(r-2) then
23:      signal Anomaly
24:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
25:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
26:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
@@
As there is a limit $n$ on the number of agents in the population, in some round $r$ there will not be enough agents to create the next $W_{r}$ in time $O(\log n)$. Our weight creation process will detect the first round $r_{max}$ in which the weight creation fails. It will be detected when an agent in state $\lambda_{r}$ or $w_{r}(i)$ for $i>0$ receives a signal to move to round $r+1$. In this case, the agent advancing to round $r+1$ will adopt state $max(r)$. This state is then propagated through the epidemic in round $r+1$ whp. The leader in state $max(r)$ when moving to round $r+2$ changes state to $Max(r)$. Then all other agents proceeding to round $r+2$ get state $Max(r)$. All these operations proceed as described above whp, calculating the value of $Max(r)$. The value of $Max(r)$ is the number assigned to the first weight that could not be created. If a follower reaches round $r+2$ with state $w_{r}(0)$, it is saved as an element of $W_{r}$ in the global part of the state. An anomaly is detected if an agent in state $max(r)$ interacts with an agent in round $r+2$ that is not in state $Max(r)$. The state $Max(r)$ occurs only in round $r+2$ and then transitions to the state endend. However, agents in the heaviest weight created are marked as top. This allows the population to know which weight is the highest, maintaining the desired number of states.
@@
WeightCreation protocol (Algorithm 3) creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$ correctly with high probability, for some value $Max(r)>\log\frac{n}{3}$, also determining this value. With negligible probability, it correctly creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$, for some $Max(r)\leq\log\frac{n}{3},$ also determining this value. Otherwise, it detects an anomaly.
@@
Let $W_{r}$ be the first weight that is not fully formed in its round. This occurs when, upon transitioning to round $r+1$, the state $\lambda_{r}$ or $w_{r}(i),$ for some $i>0,$ persists. Consequently, the state $max(r)$ emerges upon entering round $r+1$. This state is propagated whp to all agents in the population via an epidemic process during round $r+1$. When the leader advances to round $r+2$, it updates the state $max(r)$ to $Max(r)$. This causes all other agents to adopt state $Max(r)$ at the start of the round $r+2$ whp. With negligible probability, state $max(r)$ fails to reach the leader in round $r+1$. Assume also, no clock anomaly occurs. Then, the leader may advance to round $r+2$ without establishing the state $Max(r)$. This, in turn, triggers a protocol anomaly when an agent in state $max(r)$ progresses to round $r+2$. For pseudocode of Algorithm 3 see page 3. ∎
With negligible probability, the weight creation protocol may terminate prematurely. In such cases, the generated weights may be insufficient to execute the parity protocol for $X$. To ensure this does not occur in the current protocol execution, we use our majority protocol to compare the size of $X$ with the size of the largest created weight, Max(r)1Max(r)-1, after weight creation. The outcome $|X|<2^{Max(r)-1}$ guarantees that the remaining weights are sufficient to compute the parity for $X$.
3.6 Proof of Theorem 1
We start with the proof of Theorem 1, which states that there exists a population protocol for the parity problem that uses $O(\log^{3}n)$ states and silently stabilises in expected $O(\log^{3}n)$ time.
Proof of Theorem 1.

### v3

@@ 2 The Parity Problem
Recall that in the parity problem, agents are tasked with computing the parity of a given sub-population size. In this work, we present a stable population protocol that solves the parity computation problem in polylogarithmic time, see Theorem 1. We begin by presenting a fixed-state protocol that computes the parity of a given subpopulation $X$. This protocol serves as the fallback mechanism and stabilizes in $O(n\log n)$ time, both in expectation and with high probability.
@@ 2.1 Slow Stable Parity Protocol
The slow protocol builds on the leader election protocol of [5]. It uses four states: $L_{0},L_{1},F_{0},F_{1}$}, with $L_{x}$ denoting leaders and $F_{x}$ followers. Initially, all agents in subpopulation $X$ start in $L_{1}$; all others start in $F_{0}$. In the interaction of two leaders, the rule $L_{x}+L_{y}\rightarrow L_{(x+y)\bmod 2}+F_{0}$ applies, where one remains a leader (with updated parity), and the other becomes a follower. Eventually, a single leader remains, encoding the parity of $|X|$ in its state index. This is then propagated via rule $L_{x}+F_{y}\rightarrow L_{x}+F_{x}$, so all followers adopt the leader’s index.
@@
2.2 Efficient Parity Protocol
We present here an outline of the first efficient, stable, and silent parity protocol (Algorithm 1).
@@ 2.2 Efficient Parity Protocol
The balance scale idea stems from recomputation of powers of 2, as proposed in [6]. Equipped with polylogarithmic memory, we can afford the initial construction of subpopulations $W_{i}$ of cardinality $2^{i},$ for all $i=0,1,\dots,\lfloor\log\frac{n}{3}\rfloor$. These subpopulations serve as weights in balance scales implemented via the efficient majority protocol. Using subpopulation $W_{i}$, we can determine whether $|X|-|Y|\geq 2^{i}$ by computing exact majority (with tie detection) between $|X|$ and $|Y|+|W_{i}|$. This balancing mechanism enables efficient computation of parity and size of any subpopulation.
@@
There exists an efficient population protocol which solves the parity problem utilising $O(\log^{3}n)$ states and stabilises silently in expected $O(\log^{3}n)$ time.
@@ 3 Efficient Protocols with Anomaly Detection
In this section, we develop methods for constructing efficient population protocols that silently stabilise with certainty. Specifically, assuming the existence of two protocols: a slow, state-efficient protocol that stabilises silently in polynomial time with probability 1, and an efficient Monte Carlo type protocol possibly composed of subprotocols, we show how to integrate these two to obtain a protocol that is both efficient in time and guaranteed to stabilise silently.
Monte Carlo protocols typically aim to compute the desired solution with high probability, but the correctness of the solution remains uncertain after execution. Existing mechanisms for detecting efficient protocol failures and transitioning to a consistently stabilising protocol have been developed on an ad hoc basis. In contrast, we propose a novel systematic anomaly-detection mechanism that effectively integrates multiple Monte Carlo subprotocols with a reliable, though slower, always-correct solution. Specifically, this mechanism identifies efficiently scenarios necessitating a shift to a slower, always-stabilising protocol and facilitates the creation of complex efficient population protocols by sequentially combining simpler ones. Additionally, the constructed protocols are silent.
Our efficient protocols, requiring only $O(\text{poly}\log\,n)$ time and space, converge to the correct output with high probability or reliably signal potential errors through a effective anomaly-detection mechanism that enables fallback to guaranteed correctness. Such signal is treated as a failure of the efficient protocol, even if it does not always indicate an actual error. Our universal strategy is to first run an efficient Monte Carlo protocol that either stabilises with high probability or, upon detecting an anomaly, switches to a slower, always-stabilising fallback.
@@ 3.1 Clock Protocol
Central to our design is the clock protocol, which acts as a clocking device and a command counter, tracking algorithmic progress and synchronising all agents to the same point in time. In rare cases of desynchronisation, it signals an anomaly. Each main (e.g., parity) protocol is built by sequentially composing subprotocols, each spanning one or more rounds. The clock counts down $R=\text{poly}\log(n)$ rounds of the main protocol, each taking $O(\log n)$ time. To measure a round’s duration, a leader-maintained interaction counter tracks $c\log n$ interactions, triggering the transition to the next round. As in the rest of this paper, we do not focus on the best possible optimisation of the clock protocol, but on using the simplest possible mechanisms to simplify presentation.
@@ 3.2 MC+ Protocols
In this subsection, we define MC+ protocols, which compute the correct result with high probability. However, unlike regular MC protocols, they provide guarantees of correct task execution. They complete the computation after polylogarithmic time. If, for some reason, the result is incorrect, they signal an anomaly. An anomaly can also be signaled with a correct result, but its overall probability is negligible. Therefore, the absence of an anomaly guarantees the correct execution of the computation. After completing one MC+ protocol, it is possible to continue the computation using the next protocol, assuming the previous one has executed correctly. However, if the previous protocol performs the computation incorrectly, signaling an anomaly ensures a transition to a stable, but possibly slow backup. For example, the design of the leader election protocol prevents the disappearance of all leaders, which would cause the clock to stop, but it is possible that several leaders remain after its completion. The calculations of the subsequent protocols performed assuming the calculation of a single leader will not produce a sensible result, but the inevitable interaction of different leaders will lead to anomaly signaling switching the calculations to a backup.
State space.
@@
In this work, we focus on efficient population protocols, specifically Monte Carlo protocols with anomaly detection (MC+ protocols). These protocols either silently stabilise on the correct computational output whp or trigger an anomaly detection signal. To maintain order in protocol analysis, we distinguish two types of anomalies:
@@
If an efficient MC+ protocol and a polynomial-time stable protocol exist for the same problem, they can be combined to yield a protocol that is both time-efficient and stable whp. Moreover, if the polynomial-time stable protocol is silent, the resulting combined protocol is also silent.
@@
Assume that the efficient protocol completes its computations in $R$ rounds. If not successful, it reports a protocol anomaly in at most $n^{a}$ average time with negligible probability less than $n^{-a}$, for suitably chosen protocol parameters and $a>0$. In this time, by Lemma 3 the clock reports an anomaly with negligible probability less than $n^{-a}$. Assume also that the slow algorithm runs in expected time less than $n^{a}$. The expected execution time of the combined algorithms consists of the following components:
@@
Therefore, the expected execution time of the combination of efficient and slow protocol is at most $O(R\log n)+1+2=O(R\log n).$ Moreover, the combination of the protocols is stable and silent.
@@
3.3 MC+ Leader Election
The MC+ protocol for leader election proposed in this work builds upon a scheme that has been employed in several prior studies [6, 8, 13]. In this approach, the leaders perform coin flips in successive rounds. Whether an agent draws heads or tails depends on the random scheduler’s designation of that agent as either the initiator or the responder in an interaction. If, in a given round, one leader flips tails while another flips heads, the leader obtaining tails becomes a follower. Unlike previous protocols, our method associates the coin-flip outcome (heads or tails) with the round number, thereby ensuring that at least one leader always persists. Although the protocol is less efficient than several leader election protocols known from the literature, this construction reflects our emphasis on clarity and conceptual simplicity. Adapting more sophisticated existing protocols to the MC+ framework would likely entail additional technical difficulties. Furthemore, leader election does not constitute the bottleneck for the efficiency of protocols for parity or congruence.
Besides detectable clock anomalies, the only other potential failure is persistent multiple leaders, which is detected as a protocol anomaly upon interaction after leader election concludes. This detection occurs in expected time $O(n)$.
@@ 3.3 MC+ Leader Election
3.4 MC+ Exact Majority
In this section, we construct and analyse our majority computation protocol. We hypothesise that the best known protocol for majority, achieving time $O(\log n)$, could be modified into the MC+ protocol, which runs in a single or fixed number of rounds. Using such a protocol for parity computation would reduce its overall time to $O(\log^{2}n)$. Due to the complexity of this task, we have abandoned its implementation in this work and leave it as an open problem.
Our MC+ majority protocol operates over $LOG+4$ rounds and achieves a time complexity of $O(\log^{2}n)$. Throughout this construction, we assume that a unique leader has already been selected. The protocol compares two subpopulations, $X$ and $Y$, determining which is larger or whether they are of equal size. Notably, $X$ and $Y$ need not be disjoint. At the start of the first round, agents are assigned states as follows: elements exclusive to $X$ are assigned state $(1)$; elements exclusive to $Y$ state (1)(-1); elements in both $X$ and $Y$ or in neither state $(0)$. During each round $r$, an averaging operation is performed on pairs of agents in round $r$: $(a)+(b)\rightarrow\left(\lfloor\frac{a+b}{2}\rfloor\right)+\left(\lceil\frac{a+b}{2}\rceil\right)$. When transitioning to the next round, a doubling operation is applied: $(a)\rightarrow(2a)$.
@@ 3.4 MC+ Exact Majority
Given an initial configuration where each agent has an integer state $(a)$ and the state discrepancy is $\delta$, any protocol using only averaging reaches a configuration with discrepancy at most $2$ after $O(\log\delta+\log n)$ time whp.
The following theorem holds.
@@
Let bias $\Delta=|X|-|Y|$ and ${\mathcal{S}}_{r}$ be the sum of states $(a)$ in which each agent starts round $r$. Note that ${\mathcal{S}}_{1}=\Delta$. The averaging operation does not change the sum of agent states $(a)$ in round $r$. Thus as long as there are no $(x_{win})$ or $(y_{win})$ states, ${\mathcal{S}}_{r}$ is also the sum of states $(a)$ in which each agent ends round $r$. Due to doubling operation ${\mathcal{S}}_{r+1}=2{\mathcal{S}}_{r}$, as long as states $(x_{win})$ and $(y_{win})$ are not present.
@@
3.5 MC+ Weight Creation
In this section, we introduce the MC+ protocol, which creates weights that are disjoint sets of agents. Weight $W_{i}$ is a set consisting of $2^{i}$ agents. The weight creation rounds are indexed by integers $r=0,1,2,\ldots,LOG+2$. In round $r$, the protocol creates $W_{r}$, consisting of $2^{r}$ agents. If each of these agents is assigned unit mass, then $W_{r}$, composed of $2^{r}$ agents, has a total mass of $2^{r}$. At the start of the protocol, all followers are in state $\phi$, which marks free agents. At the beginning of each subsequent round $r$ the leader posessing state $\lambda$ adopts state $\lambda_{r}$, which serves as the precursor of weight $W_{r}$ carrying its mass $2^{r}$. The leader then transfers this mass to the first free agent it encounters via transition: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$, where $\lambda$ denotes a non-final leader state that carries no mass. Agents participating in the creation of $W_{r}$ have states $w_{r}(i)$, indicating that the agent possessing it belongs to $W_{r}$ and holds mass $2^{i}$ from that weight. Creating the weight is accomplished by a transition that divides the mass greater than $2^{0}=1$ between two agents: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$ when $i>0$. This process completes when all agents in $W_{r}$ have state $w_{r}(0)$ carrying mass 1.
In some round $r$, the weight $W_{r}$ will inevitably fail to unfold. This will lead to a transition to round $r+1$ with state $\lambda_{r}$ or $w_{r}(i)$ for $i>0$ present in the population. In the protocol, we mark such an event with a transition to state $m(1)$. We will propagate state $m(1)$ to the entire population via an epidemic within round $r+1$ whp. The leader moving to a new round $r+2$ in state $m(1)$ changes state to $m(2)$. Then, all followers moving to round $r+2$ in state $m(1)$ will receive state $m(2)$. We signal an anomaly if the leader does not receive state $m(1)$ during the epidemic in round $r+1$. This causes an anomaly to be signaled when a follower in state $m(1)$ attempts to transition to round $r+2$. Note that if state $m(1)$ does not appear when transitioning to round $r+1$, the weight $W_{r}$ can be stored in the global states. This is done when transitioning to round $r+2$. With negligible probability, too few weights may be created to perform the parity protocol. To detect such an event, after the weights are created, their total mass is compared to the number of elements in the set $X$ for which the parity protocol is being performed. If an insufficient number of weights are created, an anomaly is signaled.
Note that the absence of weights with masses between that of the largest created weight and $LOG$, i.e., these agent sets are empty – does not hinder the general scheme of the parity protocol.
@@ 3.5 MC+ Weight Creation
4:  for round $r\leftarrow 0$ to $LOG+2$ do
5:   upon entering round $r$ interacting with an agent in state different than $m(2)$:
6:      if $r=0$ then
7:        the leader adopts state $\lambda_{0}$ and non-leaders adopt state $\phi$.
8:      else
9:        the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $2^{r}$),
10:        a follower changes local state $w_{r-2}(0)$ to global state $w_{r-2}$ and local $w$,
11:        the leader changes state $\lambda_{r-1}$ to state $m(1)$,
12:        a follower changes state $w_{r-1}(i)$ to state $m(1)$, when $i>0$,
13:        the leader changes state $m(1)$ to state $m(2)$,
14:        a follower in state $m(1)$ signals Anomaly.
15:   upon entering round $r$ interacting with an agent in state $m(2)$:
16:      a follower adopts state $m(2)$.
17:   if interacting agents are in round $r$ then
18:       The leader transfers mass $2^{r}$ to the first free agent it encounters: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$,
19:        Distribute the mass by repeating for $i>0$: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$,
20:        Epidemic of state $m(1)$: $m(1)+\text{\emph{any state}}\rightarrow m(1)+m(1)$.
21:  Compare $\bigcup W_{k}$ and $X$ using the Majority protocol
22:    (elements of $\bigcup W_{k}$ have global states $w_{k}$)
23:  if $|\bigcup W_{k}|<|X|$ then signal Anomaly
@@
The following theorem holds.
@@
WeightCreation protocol (Algorithm 3) creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)}$ correctly with high probability, for some value $Max(r)>\log\frac{n}{6}$. With negligible probability, it correctly creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)}$, for some $Max(r)\leq\log\frac{n}{6}$. Otherwise, it detects an anomaly.
@@
Let $W_{r}$ be the first weight that is not fully formed in its round. This occurs when, upon transitioning to round $r+1$, the state $\lambda_{r}$ or $w_{r}(i),$ for some $i>0,$ persists. Consequently, the state $m(1)$ emerges upon entering round $r+1$. This state is propagated whp to all agents in the population via an epidemic process during round $r+1$. When the leader advances to round $r+2$, it updates the state $m(1)$ to $m(2)$. This causes all other agents to adopt state $m(2)$ at the start of the round $r+2$ whp. With negligible probability, state $m(1)$ fails to reach the leader in round $r+1$. Assume also, no clock anomaly occurs. Then, the leader may advance to round $r+2$ without establishing the state $m(2)$. This, in turn, triggers a protocol anomaly when a follower in state $(1)$ progresses to round $r+2$. For pseudocode of Algorithm 3 see page 3. ∎
3.6 Efficient Parity Protocol – Proof of Theorem 1
Proof.